Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 220     Entry time: 03 Jun 2023, 09:42
Author: Thomas Preston 
Group: HED 
Subject: Shift summary 

X-ray parameters: 20 keV,  352 pulses, ca 300 uJ

  • running user samples.

experiment going overall well.

X-ray performance:

  • Good - still crosstalk issues with SA3 so we have to call BKR everytime we change bunches to 30 (alignment mode), or 0 (hutch intervention).

AGIPD took damage to one module yesterday (top left) leading to a short in the high-voltage line but is still performing well with only 7/8 modules. We will investigate what the extent of the damage is after the experiment.

ELOG V3.1.4-7c3fd00